# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ 2.2 Out-of-sample estimation of Sharpe ratio
In this section, we present our proposed estimator of $SR(\mathbf{Q})$ and its main theoretical results. We need the following assumptions.
@@
The matrices $\mathbf{A}$ is deterministic. The regularization matrix $\mathbf{Q}$ is selected from some candidate set $\mathcal{Q}$, in which all the elements are independent with $\mathbf{X}$. For any $\mathbf{Q}\in\mathcal{Q}$, there exist constants $c_{\mathbf{Q}},C_{\mathbf{Q}}>0$ such that $c_{\mathbf{Q}}\leq\lambda_{\min}(\mathbf{Q})$ and $\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathop{\mathrm{op}}}}\leq C_{\mathbf{Q}}$ for any sequences $(n,p)$.
Assumption 2.5 is reasonable, as it makes no sense to apply an ill-conditioned regularization matrix $\mathbf{Q}$. The condition $\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathop{\mathrm{op}}}}\leq C_{\mathbf{Q}}$ can also hold as long as $\lambda_{\min}^{-1}({\bm{\Sigma}})$ and $\|\mathbf{Q}\|_{{\mathop{\mathrm{op}}}}$ are upper bounded by constants, which are quite mild assumptions. We then make assumptions regarding the structure of ${\bm{\Sigma}}$.
@@
${\bm{\Sigma}}$ is well scaled as $\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}\leq C$ for some constant $C>0$. Denote by $\lambda_{1}\geq\cdots\geq\lambda_{p}$ the eigenvalues of ${\bm{\Sigma}}$. One of the following cases must hold:
@@
    (Bounded spectrum) There exists $C_{1}>$ such that $0<\lambda_{1}\leq C_{1}<+\infty$ for any sequences $(n,p)$.
@@
    (Arbitrary number of diverging spikes when $c<1$) $p/n\to c<1$ and we allow arbitrary number of top eigenvalues goes to infinity. It is assumed that there exists constant $c_{1}>0$ such that $\lambda_{p}\geq c_{1}$.
@@
    (Fixed number of diverging spikes when $c\geq 1$) $p/n\to c\geq 1$ and we let the number of diverging spikes be $K$, $K$ is fixed and $\lambda_{1}/\lambda_{K}\leq C_{\lambda}$ for some constant $C_{\lambda}>0$. Moreover, it is assumed that there exists a constant $c_{1}>0$ such that $\lambda_{p}\geq c_{1}$.
Assumption 2.6 includes three cases: the first case is where ${\bm{\Sigma}}$ has bounded operator norm, and the second/third case is where ${\bm{\Sigma}}$ has spikes that tend to infinity. This encompasses a wide range of structures of the covariance matrix ${\bm{\Sigma}}$. Specifically, we allow the number of spikes to tend to infinity when $c<1$, with flexibility in the rate at which they tend to infinity. While the condition becomes slightly stronger for $c\geq 1$, it remains mild in practical settings as we only need to make sure we have a fixed number of diverging eigenvalues. The main results are given in the following theorem.
@@
If additionally $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded [^1] The condition can be further relaxed to ${\mathrm{tr}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}$ is bounded, where $s_{0}>0$ is defined by $s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}$ in Assumption 3.1 below., then $SR(\mathbf{Q})$ is almost surely bounded and
@@
The proof of Theorem 2.7 is given in Appendix D in the supplementary material. In the proof, we present several extensions of Lemma C.1, which is a key result from random matrix theory. The first challenge in the extension arises from the relaxation of bounded $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ in Lemma C.1. When spikes exist and $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is no longer bounded, we cannot apply Lemma C.1 and have to extend Lemma C.1 to Lemmas C.6 and C.7, and prove that $z\to 0$ and $n\to+\infty$ can be interchanged. The second challenge arises from obtaining ratio consistency as well as difference consistency. As we see, ratio consistency has the advantage of circumventing additional scaling condition (i.e. bounded $\|\mathbf{A}\|_{{\mathrm{tr}}}$). To obtain ratio consistency, we establish Lemma C.9 even when $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is unbounded.
Theorem 2.7 establishes an estimator for $SR(\mathbf{Q})$ with almost sure convergence guarantee in a wide range of scenarios. More importantly, it provides a tractable framework for portfolio managers to determine the best regularization matrix $\mathbf{Q}$ over some given candidate set $\mathcal{Q}$. Consider a candidate set $\mathcal{Q}$ from which we want to select $\mathbf{Q}\in\mathcal{Q}$ in order to generate the maximum out-of-sample Sharpe ratio. It holds from Theorem 2.7 that $\widehat{SR}(\mathbf{Q})/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, hence we can simply maximize $\widehat{SR}(\mathbf{Q})$ over $\mathcal{Q}$ to obtain the optimal $\mathbf{Q}$. However, it is also important to note that optimizing the entire matrix $\mathbf{Q}$ without the constraint $\mathbf{Q}\in\mathcal{Q}$ to obtain the maximum value of $SR(\mathbf{Q})$ is not feasible. In other words, if we define $\widehat{\mathbf{Q}}=\arg\max_{\mathbf{Q}}\widehat{SR}(\mathbf{Q})$, we cannot guarantee that $SR(\widehat{\mathbf{Q}})$ will also be large, that is, $\widehat{SR}(\widehat{\mathbf{Q}})$ may not align well with $SR(\widehat{\mathbf{Q}})$. In our theorem, $\mathbf{Q}$ is independent of $\mathbf{X}$, but here $\widehat{\mathbf{Q}}$ is derived from the observed data matrix $\mathbf{X}$. Nonetheless, the estimator is practically useful to maximize $SR(\mathbf{Q})$ within a suitable candidate set $\mathcal{Q}$ without access to the true ${\bm{\Sigma}}$. For instance, we can let $\mathcal{Q}$ be the set that exists countably many deterministic matrices. The final important comment is that, with well designed $\mathbf{Q}$, we show in Appendix F that the portfolio optimization can indeed achieve the maximum population SR $=\sqrt{{\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}}$.
@@ 3.1 Efficient frontier with no risk-free asset
$\hfil~{}&\mathbf{g}=D^{-1}\big{[}B(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}-A(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}\big{]},~{}\mathbf{h}=D^{-1}\big{[}C(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}-A(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}\big{]},\hfil~{}\\ \hfil~{}&A=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}B=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r},~{}C={\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}D=BC-A^{2}.\hfil~{}\end{array}$ (3.5)
@@
which should be viewed as the curve of $(\sigma_{0},\mu_{0})$ as we change the target return $\mu_{0}$. Our objective is to estimate this curve represented by (3.6) for a given regularization matrix $\mathbf{Q}$.
@@ 3.2 Out-of-sample estimation of frontier variances
In this section, we present the main results on estimating the efficient frontier. We give an additional technical assumption on $\mathbf{r}$, assuming that the vectors $\mathbf{r}$ and ${\bm{1}}$ do not coincide, which makes $\mu_{0}=1$ so that we do not have a well-posed problem.
@@
The proof of Theorem 3.2 can be found in Appendix E in the supplementary material. It is worth noting that when the target return $\mu_{0}$ is given, estimating the volatility naturally leads to an estimator of Sharpe ratio. The optimal true Sharpe ratio $SR(\mathbf{Q})$ without the risk-free asset equals $(\mu_{0}-r_{0})/\sigma_{0}$, and Theorem 3.2 implies that the estimator $\widehat{SR}(\mathbf{Q})=(\mu_{0}-r_{0})/\widehat{\sigma}$ is ratio-consistent without additional scaling condition, and is also difference-consistent when $\mathcal{A}_{rr}$ is bounded. In Theorem 2.7, if we set $\mathbf{A}=\mathbf{r}\mathbf{r}^{\top}$, the relaxed condition that ${\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}$ is bounded (see the footnote for Theorem 2.7) for the difference-consistency of Sharpe ratio is exactly the same as the condition that $\mathcal{A}_{rr}$ is bounded in Theorem 3.2.
@@
Albeit the conditions are similar, we hope to point out new challenges in proving Theorem 3.2. The first challenge arises from the presence of distant diverging spikes in the covariance matrix ${\bm{\Sigma}}$ in Case 2 and Case 3 of Assumption 2.6. It is possible that the vectors $\mathbf{r}$ and ${\bm{1}}$ lie within the space spanned by the eigenvectors of these spikes. Note that in Section 2, due to the existence of the risk-free asset, we directly work with the exccess return ${\bm{\mu}}$ without necessity to discuss the relationship between $\mathbf{r}$ and ${\bm{1}}$. For instance, we consider the following 4 concrete cases with the assumption $\mathbf{r}=a\cdot{\bm{1}}+\mathbf{r}^{\perp}$ where $\mathbf{r}^{\perp}\perp{\bm{1}}$, and the distant spikes in ${\bm{\Sigma}}/(1+s_{0})+\mathbf{Q}$ are all in the order of $p$.
  • Case 1: ${\bm{1}}$ and $\mathbf{r}^{\perp}$ are both orthogonal to the space of distant spikes, or no spikes.
  • Case 2: ${\bm{1}}$ belongs to, and $\mathbf{r}^{\perp}$ is orthogonal to the space of distant spikes.
  • Case 3: $\mathbf{r}^{\perp}$ belongs to, and ${\bm{1}}$ is orthogonal to the space of distant spikes.
  • Case 4: ${\bm{1}}$ and $\mathbf{r}^{\perp}$ both belong to the space of distant spikes.
Table 1: Different rates of $\mathcal{A}_{rr}$ and $\mathcal{A}_{11}$ for specific cases.
Case 1 Case 2 Case 3 Case 4
$\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$ $\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$ $\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$ $\|\mathbf{r}^{\perp}\|_{2}=1$ $\|\mathbf{r}^{\perp}\|_{2}=\sqrt{p}$
Order of $\mathcal{A}_{rr}$ $\Theta(a^{2}p+1)$ $\Theta(a^{2}p+p)$ $\Theta(a^{2}+1)$ $\Theta(a^{2}+p)$ $\Theta\big{(}a^{2}p+\frac{1}{p})$ $\Theta(a^{2}p+1)$ $\Theta\big{(}a^{2}+\frac{1}{p}\big{)}$ $\Theta(a^{2}+1)$
Order of $\mathcal{A}_{11}$ $\Theta(p)$ $\Theta(p)$ $\Theta(1)$ $\Theta(1)$ $\Theta(p)$ $\Theta(p)$ $\Theta(1)$ $\Theta(1)$
$\frac{\widehat{\sigma}^{2}}{\sigma_{0}^{2}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
$\frac{\mu_{0}-r_{0}}{\sigma_{0}}-\frac{\mu_{0}-r_{0}}{\widehat{\sigma}}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$ $a=0$ $\surd$ - $\surd$ - $\surd$ $\surd$ $\surd$ $\surd$
$a=\frac{1}{\sqrt{p}}$ $\surd$ - $\surd$ - $\surd$ $\surd$ $\surd$ $\surd$
$a=1$ $\surd$ - - - $\surd$ $\surd$
Range of $\mu_{0}$ $O(\sqrt{a^{2}p+1})$ $O(\sqrt{a^{2}p+p})$ $O(\sqrt{a^{2}+1})$ $O(\sqrt{a^{2}+p})$ $O\big{(}\sqrt{a^{2}p+\frac{1}{{p}}}\big{)}$ $O(\sqrt{a^{2}p+1})$ $O\big{(}\sqrt{a^{2}+\frac{1}{{p}}}\big{)}$ $O(\sqrt{a^{2}+1})$
for $\widehat{\sigma}^{2}-\sigma_{0}^{2}~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0$
The symbol $\surd$” indicates that the convergence is proven to hold.
Table 1 gives the different rates of $\mathcal{A}_{rr}$ and $\mathcal{A}_{11}$ for the above 4 cases, all of which are covered in Theorem 3.2. From Table 1, it is obvious that the ratio-consistency always holds without additional conditions. The Sharpe difference-consistency holds only when $\mathcal{A}_{rr}$ is bounded, which becomes harder to be satisfied if we bet the market factor ${\bm{1}}$ in the mean return. This partially explains why portfolio managers eagerly pursue to be market neutral, that is $a=0$. Furthermore, the target return $\mu_{0}$ needs to be set within a proper range in order for the corresponding volatility to be estimated accurately in terms of the absolute error. We can usually allow $\mu_{0}$ to be on the constant order, except in Case 3 and Case 4 when $\|\mathbf{r}^{\perp}\|=1$ and $a=0$, which represents a very low signal level in $\mathbf{r}$.
Handling general $\mathcal{A}_{rr},\mathcal{A}_{11},\mathbf{r}$ and $\mu_{0}$ with arbitrary rates poses great challenge in the analysis. In the proof of Theorem 3.2, we derive the orders of all the relevant quantities in different cases in terms of $\mathcal{A}_{rr}$, $\mathcal{A}_{11}$ and $\mu_{0}$, without specifying their specific statistical rates. Additionally, we need to divide the proof into two separate cases where $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ and $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, where distinct analyses are required.
In addition to the above order-related challenge mentioned, the expression of $\sigma_{0}^{2}$ presents another obstacle. While $\sigma_{0}^{2}$ can be expressed as ${\bm{\xi}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\xi}}$ for some vector ${\bm{\xi}}$, similar to (2.3) in Section 2, it is important to note that ${\bm{\xi}}$ depends on $\widehat{\bm{\Sigma}}$. To tackle this nontrivial dependency, we need to carefully substitute ${\bm{\xi}}$ with a deterministic counterpart and make sure this does not affect the almost sure convergence. More proof details can be found in Appendix E in the supplementary material.

### v3

@@ 2.2 Out-of-sample estimation of Sharpe ratio
In this section, without any information of ${\bm{\Sigma}}$, we propose a way to effectively estimate the Sharpe ratio $SR(\mathbf{Q})$ in (2.4) with in-sample data. We need the following assumptions.
@@
The matrices $\mathbf{A}$ is deterministic. The positive definite regularization matrix $\mathbf{Q}\in\mathcal{Q}$, independent of $\mathbf{X}$, satisfies the following: there exist constants $c_{1},c_{2}\geq 0$, with $c_{1}^{2}+c_{2}^{2}>0$, and $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$ such that $\mathbf{Q}=c_{1}\mathbf{Q}_{1}+c_{2}\mathbf{Q}_{2}$. The matrices $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$ satisfy $c^{\prime}\mathbf{I}\leq\mathbf{Q}_{1}\leq C^{\prime}\mathbf{I}$ and $c^{\prime}\mathbf{I}\leq{\bm{\Sigma}}^{-1/2}\mathbf{Q}_{2}{\bm{\Sigma}}^{-1/2}\leq C^{\prime}\mathbf{I}$ for some constants $c^{\prime},C^{\prime}>0$, for any sequence $(n,p)$. In addition, we allow the case $\mathbf{Q}={\bm{0}}$ when $c<1$.
Assumption 2.5 says that it makes no sense to apply an ill-conditioned regularization matrix $\mathbf{Q}$. For instance, we can choose $\mathbf{Q}$ with bounded eigenvalues by setting $c_{1}=1$, $c_{2}=0$, $\mathbf{Q}_{1}=\mathbf{Q}$. $\mathbf{Q}$ can also be proportional to ${\bm{\Sigma}}$, even when $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is unbounded, by setting $c_{1}=0$, $c_{2}=1$, and $\mathbf{Q}_{2}={\bm{\Sigma}}$. We also allow no regularization $\mathbf{Q}={\bm{0}}$ when $c<1$.
@@
${\bm{\Sigma}}$ is well scaled as $\|{\bm{\Sigma}}/p\|_{{\mathrm{tr}}}\leq C$ for some constant $C>0$. Denote by $\lambda_{1}\geq\cdots\geq\lambda_{p}$ the eigenvalues of ${\bm{\Sigma}}$. It is assumed that $\lambda_{p}\geq c_{1}$ for some constant $c_{1}>0$. Moreover, one of the following cases must hold:
@@
    (Bounded spectrum) There exists $C_{1}>0$ such that $\lambda_{1}\leq C_{1}$.
@@
    (Arbitrary number of diverging spikes when $c<1$) $p/n\to c<1$ and we allow arbitrary number of top eigenvalues to go to infinity.
@@
    (Fixed number of diverging spikes when $c\geq 1$) $p/n\to c\geq 1$ and we let the number of diverging spikes be $K$, $K$ is fixed and $\lambda_{1}\leq C\lambda_{K}^{2}$ for some constant $C>0$.
Assumption 2.6 is a very mild assumption and encompasses a wide range of structures of the covariance ${\bm{\Sigma}}$. Case 1 assumes that ${\bm{\Sigma}}$ has bounded operator norm, which was typically adopted by RMT, such as Li et al. (2022) and Bodnar et al. (2024). Cases 2 and 3 correspond to the factor model structure, where ${\bm{\Sigma}}$ has spiked eigenvalues that tend to infinity. Specifically, Bai (2003) and Fan et al. (2013) considered a factor model with fixed $K$ number of strong or pervasive factors with $\lambda_{1}\asymp\lambda_{K}\asymp p$, which falls in Case 2 or 3 depending on the dimensionality regime $p<n$ or $p>n$. Onatski (2012) and Wang & Fan (2017) considered semi-strong or weak factors, that is, $\lambda_{k}\asymp p^{\theta_{k}}$ for some $\theta_{k}\in(0,1)$ and $\theta_{k}$’s may not be the same. Our Case 2 can allow arbitrary $\theta_{k}$ when $c<1$ and our Case 3 can allow flexibility in setting $\theta_{k}$ to some degree ($\theta_{K}\leq\theta_{1}\leq 2\theta_{K}$). In addition, Assumption 2.6 actually allows unbounded diagonal elements of ${\bm{\Sigma}}$. Ao et al. (2019) and the above factor model literature assume that the diagonal elements ${\bm{\Sigma}}_{i,i}$ are uniformly bounded. Here, we only require that the average $\sum_{i=1}^{p}{\bm{\Sigma}}_{i,i}/p$ remains bounded.
@@
If additionally $\|\mathbf{A}\|_{{\mathrm{tr}}}$ is bounded [^2] The condition can be further relaxed to bounded ${\mathrm{tr}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}\mathbf{A}$, where $s_{0}>0$ is defined by $s_{0}=\frac{c}{p}{\mathrm{tr}}~{}{\bm{\Sigma}}\big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\big{)}^{-1}$ in Assumption 3.1 below., then $SR(\mathbf{Q})$ is almost surely bounded and
@@
The proof of Theorem 2.7 is given in Appendix D in the supplementary material. The key in our proof is to extend a classical result (see Lemma C.1) in random matrix theory. The first challenge in the extension arises from the relaxation of bounded $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ in Lemma C.1. When spikes exist and $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is no longer bounded, we cannot apply Lemma C.1 and have to extend Lemma C.1 to Lemmas C.6 and C.7, and prove that $z\to 0$ and $n\to+\infty$ can be interchanged. The second challenge arises from obtaining ratio consistency as well as difference consistency. As we see, ratio consistency has the advantage of circumventing additional scaling condition (i.e. bounded $\|\mathbf{A}\|_{{\mathrm{tr}}}$). To obtain ratio consistency, we establish Lemma C.9 even when $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ is unbounded.
Theorem 2.7 establishes an estimator for $SR(\mathbf{Q})$ with almost sure convergence guarantee in a wide range of scenarios. More importantly, it provides a tractable framework for portfolio managers to determine the best regularization matrix $\mathbf{Q}$ over some given candidate set $\mathcal{Q}$. Consider a candidate set $\mathcal{Q}$ from which we want to select $\mathbf{Q}\in\mathcal{Q}$ in order to generate the maximal out-of-sample Sharpe ratio. It holds from Theorem 2.7 that $\widehat{SR}(\mathbf{Q})/SR(\mathbf{Q})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1$, hence we can simply maximize $\widehat{SR}(\mathbf{Q})$ over $\mathcal{Q}$ to obtain the optimal $\mathbf{Q}$. However, it is also important to note that if we define $\widehat{\mathbf{Q}}=\arg\max_{\mathbf{Q}}\widehat{SR}(\mathbf{Q})$ without the constraint $\mathbf{Q}\in\mathcal{Q}$, $\widehat{SR}(\widehat{\mathbf{Q}})$ may not align well with $SR(\widehat{\mathbf{Q}})$. This misalignment arises because optimizing $\widehat{SR}(\mathbf{Q})$ over the entire $\mathbf{Q}$ can overfit to the in-sample observed data, potentially inflating $\widehat{SR}(\widehat{\mathbf{Q}})$ and creating discrepancy between $\widehat{SR}(\widehat{\mathbf{Q}})$ and $SR(\widehat{\mathbf{Q}})$. Readers may refer to Appendix A.3 for numerical results demonstrating this discrepancy. Nonetheless, optimizing $\mathbf{Q}$ over a finite deterministic candidate set $\mathcal{Q}$, or over $\mathcal{Q}$ parameterized by a finite number of flexible parameters (see Proposition 2.8 below), can be practically useful.
Proposition 2.8.
Given the conditions of Theorem 2.7, define $\widehat{\mathbf{Q}}=\arg\max_{\mathbf{Q}\in\mathcal{Q}}\widehat{SR}(\mathbf{Q})$. Suppose $\mathcal{Q}$ has finite degree of freedom, under assumptions in Appendix I, it holds that
$\widehat{SR}(\widehat{\mathbf{Q}})/SR(\widehat{\mathbf{Q}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}1\quad\text{and}\quad\widehat{SR}(\widehat{\mathbf{Q}})-SR(\widehat{\mathbf{Q}})~{}{\stackrel{{\scriptstyle a.s}}{{\rightarrow}}}~{}0.$
Following Theorem 2.7, a natural question is whether $SR(\mathbf{Q})$ can achieve or approximate the theoretical maximal Sharpe ratio $SR_{\max}$, where $SR_{\max}=\sqrt{{\bm{\mu}}^{\top}{\bm{\Sigma}}^{-1}{\bm{\mu}}}$. In Appendix F, we show that for any given $\varepsilon>0$, there always exists $\mathbf{Q}$ such that $1-\varepsilon\leq SR(\mathbf{Q})/SR_{\max}\leq 1$ with probability one as $n\to+\infty$. The existence of $\mathbf{Q}$ is proved by setting $\mathbf{Q}=q{\bm{\Sigma}}$ for some sufficiently large $q>0$. When the covariance follows from a factor model, that is ${\bm{\Sigma}}=\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}$ with the factor loading matrix $\mathbf{B}\in\mathbb{R}^{p\times K}$ and the residual covariance $\mathbf{D}$, which is typically assumed as a diagonal matrix with eigenvalues bounded away from 00 and +$\infty$+ ∞, then the existence of $\mathbf{Q}$ can also be proved by setting $\mathbf{Q}$ proportional to $\mathbf{D}$. An experienced and skillful portfolio manager may set an initial regularization matrix $\mathbf{Q}_{0}$ as closely aligned (or proportional) to the diagonal matrix of residual variances as possible, say using some extra data to estimate the residual variances. But sometimes it may be inevitable to suffer a loss in $SR(\mathbf{Q})$ from the maximal $SR_{\max}$ practically, as we only have partial information on ${\bm{\Sigma}}$. Nonetheless, with any given $\mathbf{Q}_{0}$, by Theorem 2.7, we can further explore $\mathbf{Q}=q\cdot\mathbf{Q}_{0}$ and fine-tune the scale (or shrinkage level) to achieve the best out-of-sample SR within this class of $\mathbf{Q}$’s.
@@
Theorem 2.7 analyzes the out-of-sample Sharpe ratio under the assumption that ${\bm{\mu}}$ is known. When ${\bm{\mu}}$ is unknown and estimated by the sample mean $\widehat{{\bm{\mu}}}$, the out-of-sample Sharpe ratio is given by $\frac{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q}){\bm{\mu}}}{\sqrt{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}}}$. Under a slightly stronger assumption, where the data follow the Gaussian distribution, we also propose a corresponding estimator for the out-of-sample Sharpe ratio $\widehat{SR}(\mathbf{Q})=\frac{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}-\frac{{\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}}{n-{\mathrm{tr}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}}}{\sqrt{\bm{\widehat{\mu}}^{\top}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{{\bm{\Sigma}}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}}}\cdot\big{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\big{)}$. Detailed simulations, real data results, and theoretical justification on this estimator can be found in Appendix A, B and G, respectively.
@@ 3.1 Efficient frontier with no risk-free asset
$\hfil~{}&\mathbf{g}=\frac{B}{D}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}}-\frac{A}{D}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r},~{}\mathbf{h}=\frac{C}{D}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r}-\frac{A}{D}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},\hfil~{}\\ \hfil~{}&A=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}B=\mathbf{r}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\mathbf{r},~{}C={\bm{1}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{1}},~{}D=BC-A^{2}.\hfil~{}\end{array}$ (3.5)
@@
which should be viewed as a curve of $(\sigma_{0},\mu_{0})$ as we change the target return $\mu_{0}$. Our objective is to estimate this curve represented by (3.6) for a given regularization matrix $\mathbf{Q}$.
@@ 3.2 Out-of-sample estimation of frontier variances
In this section, we present the main results on estimating the efficient frontier. We give an additional technical assumption on $\mathbf{r}$, assuming that the vectors $\mathbf{r}$ and ${\bm{1}}$ do not coincide. Otherwise, we have $\mu_{0}=1$ and we do not have a well-posed efficient frontier.
@@
The proof of Theorem 3.2 can be found in Appendix E in the supplementary material. Notice that when the target return $\mu_{0}$ is given, estimating the volatility naturally leads to an estimator of Sharpe ratio. The optimal true $SR(\mathbf{Q})$ without the risk-free asset equals $(\mu_{0}-r_{0})/\sigma_{0}$, and Theorem 3.2 implies that the estimator $\widehat{SR}(\mathbf{Q})=(\mu_{0}-r_{0})/\widehat{\sigma}$ is ratio-consistent without additional scaling condition, and is also difference-consistent when $\mathcal{A}_{rr}$ is bounded. In Theorem 2.7, if we set $\mathbf{A}=\mathbf{r}\mathbf{r}^{\top}$, the relaxed condition that ${\mathrm{tr}}\Big{(}\frac{{\bm{\Sigma}}}{1+s_{0}}+\mathbf{Q}\Big{)}^{-1}\mathbf{A}$ is bounded (see the footnote for Theorem 2.7) for the difference consistency of Sharpe ratio is exactly the same as the condition that $\mathcal{A}_{rr}$ is bounded in Theorem 3.2.
@@
Albeit the conditions are similar, we hope to point out new challenges in proving Theorem 3.2. The first challenge arises from the presence of distant diverging spikes in the covariance matrix ${\bm{\Sigma}}$ in Case 2 and Case 3 of Assumption 2.6. It is possible that the vectors $\mathbf{r}$ and ${\bm{1}}$ lie within the space spanned by the eigenvectors of these spikes. Note that in Section 2, due to the existence of the risk-free asset, we directly work with the exccess return ${\bm{\mu}}$ without necessity to discuss the relationship between $\mathbf{r}$ and ${\bm{1}}$. But in the proof of Theorem 3.2, we need to handle general $\mathcal{A}_{rr},\mathcal{A}_{11},\mu_{0}$ without specifying their statistical rates and without specifying a detailed relationship between $\mathbf{r}$ and ${\bm{1}}$. Readers may refer to Appendix H for an example illustrating the values of $\mathcal{A}_{rr}$, $\mathcal{A}_{11}$, $\mathcal{A}_{r1}$, the possible range of $\mu_{0}$, and how $\mathcal{A}_{r1}^{2}/(\mathcal{A}_{11}\mathcal{A}_{rr})\leq\rho<1$ can be satisfied under mild conditions. Moreover, we need to divide the proof into the cases $\mu_{0}\leq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$ and $\mu_{0}\geq C\sqrt{\mathcal{A}_{rr}/\mathcal{A}_{11}}$, where distinct analyses are required. In addition to the above order-related challenge, the expression of $\sigma_{0}^{2}$ presents another obstacle. While $\sigma_{0}^{2}$ can be expressed as ${\bm{\xi}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\xi}}$ for some vector ${\bm{\xi}}$, similar to (2.3) in Section 2, it is important to note that ${\bm{\xi}}$ depends on $\widehat{\bm{\Sigma}}$. To tackle this nontrivial dependency, we need to carefully substitute ${\bm{\xi}}$ with a deterministic counterpart and make sure this does not affect the almost sure convergence. More proof details can be found in Appendix E in the supplementary material.